What is the x intercept for y=2x^2-x-3?
I got -1.1 not sure if that is right.
Anne

Answer by J2R2R(94)

2x^2 - x - 3 = (2x - 3)(x + 1) which must equal 0 on the x-axis.
So, either 2x – 3 = 0 gives x = 1.5 or x + 1 = 0 gives x= -1.
Therefore the intercepts of the x-axis where y = 0 is where x = -1 and x = 1.5